OpenStudy (anonymous):

2 angles are supplementary. one angle measures 12 degrees more than the other. find the measures of these angles.

OpenStudy (anonymous):

One angle is x Other angle is x +12 x +x +12 = 180 Solve for x 2x +12 = 180 2x = 168 x = 84 x +12 = 84 +12 = 96